OpenStudy (anonymous):

Which of the following are geometric sequences? Check all that apply. A. -9, -3, -1, -1/3, -1/9, -1/27 B. 5, 10, 15, 20, 25 C. 1, 1, 2, 3, 5, 8, 13 D. 10, 5, 2.5, 1.25, 0.625, 0.3125

OpenStudy (callisto):

A, D

OpenStudy (callisto):

A: r = 1/3, D: r=1/2 B is an AS and C is the fibonacci sequence

Directrix (directrix):

A.) -9, -3, -1, -1/3, -1/9, -1/27 D.) 10, 5, 2.5, 1.25, 0.625, 0.3125